a 1.75 kg stone falls from the top of a cliff and strikes the ground at a speed of 58.8
m/s. what is the height of the cliff?
1
Expert's answer
2020-01-29T13:58:09-0500

We can find the height of the cliff from the Law of Conservation of Energy:


"PE = KE,""mgh = \dfrac{1}{2}mv^2,""h = \dfrac{v^2}{2g} = \dfrac{(58.8 \dfrac{m}{s})^2}{2 \cdot 9.8 \dfrac{m}{s^2}} = 176.4 m."

Answer:

"h = 176.4 m."